# Structured methods for parameter inference and uncertainty quantification for mechanistic models in the life sciences

### v1

@@ Abstract
Parameter inference and uncertainty quantification are important steps when relating mathematical models to real-world observations, and when estimating uncertainty in model predictions. However, methods for doing this can be computationally expensive, particularly when the number of unknown model parameters is large. The aim of this study is to develop and test an efficient profile likelihood-based method, which takes advantage of the structure of the mathematical model being used. We do this by identifying specific parameters that affect model output in a known way, such as a linear scaling. We illustrate the method by applying it to three caricature models from different areas of the life sciences: (i) a predator-prey model from ecology; (ii) a compartment-based epidemic model from health sciences; and, (iii) an advection-diffusion-reaction model describing transport of dissolved solutes from environmental science. We show that the new method produces results of comparable accuracy to existing profile likelihood methods, but with substantially fewer evaluations of the forward model. We conclude that our method could provide a much more efficient approach to parameter inference for models where a structure approach is feasible. Code to apply the new method to user-supplied models and data is provided via a publicly accessible repository.
@@ Introduction
In 2014, Hines et al. [6] reviewed the application of Bayesian Markov chain Monte Carlo (MCMC) sampling methods for parameter estimation and parameter identifiability for a range of ordinary differential equation (ODE)-based models of chemical and biochemical networks. Their results indicated that parameter non-identifiability can be detected through MCMC chains failing to converge. Similar observations were made by Siekmann et al. [7] for a range of continuous-time Markov chain models used in the study of cardiac electrophysiology. In 2020, Simpson et al. [8] compared MCMC sampling with a profile likelihood-based method for parameter inference and parameter identifiability for a range of nonlinear partial differential equation (PDE)-based models used to interrogate cell biology experiments. The authors studied identifiable and non-identifiable problems and found that, while both the MCMC and profile likelihood-based approaches gave similar results for identifiable models, only the profile likelihood approach provided mechanistic insight for non-identifiable problems. They also found that profile likelihood was approximately an order of magnitude faster to implement than MCMC sampling, regardless of whether the problem was identifiable or not. The speed-up in computation is a consequence of the fact that it is often faster to use numerical optimisation methods compared to sampling methods. Raue et al. [9] suggested the sequential use of profile likelihood to constrain prior distributions before applying MCMC sampling in the face of non-identifiability.
While profile likelihood has long been used to assess parameter identifiability and parameter estimation [10, 11, 12, 13, 14], Simpson and Maclaren [15] recently presented a profile likelihood-based workflow covering identifiability, estimation and prediction. The workflow uses computationally efficient optimisation-based methods to estimate the maximum likelihood, and then explores the curvature of the likelihood through a series of univariate profile likelihood functions that target one parameter at a time. This workflow then propagates uncertainty in parameter estimates into model predictions to provide insight into how variability in data leads to uncertainty in model predictions. While the initial presentation of the profile likelihood-based workflow focused on deterministic ODE and PDE models, the same ideas can be applied to stochastic models whenever a surrogate likelihood function is available [1].
@@
As a motivating example, [16] described an epidemiological model that was fitted to data in real-time and used to provide policy advice during the Covid-19 pandemic. The model consisted of a system of several thousand ODEs, and model fitting and uncertainty quantification was done using an approximate Bayesian computation method targeting 11 unknown parameters. This resulted in a computationally expensive problem with a high-dimensional parameter space. Two of the fitted parameters were multiplicative factors on the proportion of model infections that led to hospitalisation and death respectively in each age group and susceptibility class. Adjusting either or both of these parameters while holding the other parameters fixed would linearly scale the time series for expected daily hospital admissions and deaths output by the model. Thus, for a given combination of the other nine fitted parameters, it should be possible to find the optimal values for these two parameters without a costly re-evaluation of the full model.
Here, we propose an new approach, which we term structured inference, that exploits the known scaling relationship between certain parameters and the model solution. This approach recasts an $N$-dimensional optimisation problem as a nested pair of lower-dimensional problems, effectively reducing the dimensionality of the search space. We illustrate our method using canonical toy models representing three cases studies drawn from different areas of the life and physical sciences: ecological species interactions; epidemiological dynamics; and pollutant transport and deposition. In each case study, we use the model structure to identify parameters that have a known scaling effect on model output. We then show that implementing a structured inference approach results in solutions with very similar accuracy, but substantially fewer calls to the forward model solver. This suggests that our method could potentially offer significant reduction in computation time when applied to more complex models, which are computationally expensive to solve. Alongside this article, we provide fully documented code for implementing our method, with instructions for how it can be applied to user-supplied models and data.
@@ Methods
$y(\theta_{\mathrm{outer}},\theta_{\mathrm{inner}})=F\left(y(\theta_{\mathrm{outer}},\theta_{\mathrm{inner}}^{\mathrm{ref}})\right)$ (2)
for some known function $F$, which we assume is relatively cheap to compute compared to solving the full forward model for $y(\theta)$. The simplest example is where the model solution is directly proportional $\theta_{\mathrm{inner}}$, in which case this transformation is a simple linear scaling:
@@
For the structured inference method, we set $\theta_{\mathrm{inner}}=p_{\mathrm{obs}}$, noting that if ${\bf y}_{\mathrm{ref}}$ is the model solution with $\theta_{\mathrm{outer}}=\theta_{\mathrm{outer}}^{\mathrm{ref}}$ and $p_{\mathrm{obs}}=1$, then the solution for a arbitrary value of $p_{\mathrm{obs}}\in[0,1]$ is simply ${\bf y}=p_{\mathrm{obs}}{\bf y}_{\mathrm{ref}}$. Thus, under the structured method, the ordinary differential equation model (7)–(8) only ever needs to be solved with $p_{\mathrm{obs}}=1$. When evaluation of the model likelihood requires the expected value ${\bf y}$ of observed data for any other value of $p_{\mathrm{obs}}$, this is obtained via this simple linear scaling.
SEIR model: Epidemiology
Our second case study is a SEIR compartment model for an epidemic in a closed population of size $N$ [22]. We assume that the population at time $t$ is divided into compartments representing the number of people who are susceptible ($S$), exposed ($E$), infectious ($I$), and recovered ($R$). To model waning immunity, we subdivide the recovered compartment into two compartments $R_{1}$ and $R_{2}$ and assume that individuals transition from $R_{1}$ to $R_{2}$ and subsequently back to $S$. This is described by the following system of ordinary differential equations:
@@
We assume that, on average, a fixed proportion $p_{\mathrm{obs}}$ of infections are observed. This could represent epidemiological surveillance data, for example the number of notified cases under the assumption that surveillance effort is steady over time. Alternatively it could represent a particular clinical outcome, such as hospital admission. Unknown case ascertainment or uncertain clinical severity are common, particularly for new or emerging infectious diseases. Hence, joint inference of this parameter with parameters of the SEIR model, in situations where these are identifiable, is an important task in outbreak modelling [16, 23, 24].
@@
We assume that the number of observed cases $\hat{X}_{t}$ on day $t$ is drawn from a negative binomial distribution with mean $y_{t}$ and dispersion factor $k$:
@@
Similar to the predator–prey model, we set $\theta_{\mathrm{inner}}=p_{\mathrm{obs}}$, noting that if ${\bf y}_{\mathrm{ref}}$ is the model solution with $\theta_{\mathrm{outer}}=\theta_{\mathrm{outer}}^{\mathrm{ref}}$ and $p_{\mathrm{obs}}=1$, then the solution for a arbitrary value of $p_{\mathrm{obs}}\in[0,1]$ is simply ${\bf y}=p_{\mathrm{obs}}{\bf y}_{\mathrm{ref}}$.
@@
$u(x,t)$ $=$ $\dfrac{U_{0}}{2}\left[\textrm{erfc}\left(\dfrac{x-tv^{*}}{\sqrt{4tD^{*}}}\right)+\textrm{exp}\left(\dfrac{xv^{*}}{D^{*}}\right)\textrm{erfc}\left(\dfrac{x+tv^{*}}{\sqrt{4tD^{*}}}\right)\right],$ (21)
@@
For the structured method, we exploit a symmetry in the analytical solution in Eqs.(21)–(22), namely that
$u(x,t;R)=u\left(x,(R_{0}/R)t;R_{0}\right).$ (24)
We therefore set $\theta_{\mathrm{inner}}=R$ and note that if $u^{\mathrm{ref}}$ is the model solution when $R=1$, then the solution for an arbitrary value of $R>1$ is $u(x,t)=u^{\mathrm{ref}}(x,t/R)$. The transformation required to obtain the model solution is more complicated than the simple linear scaling that existed in the first two case studies, and requires evaluating the solution for $u(x,t)$ at a different value of $t$. However, since a numerical scheme for an advection–diffusion PDE will typically generate solution values at a series of closely spaced time points, this information will generally be available from the original call to the numerical solver with $R=1$, and will not require costly re-evaluation of the numerical solution scheme. To emulate this situation, we evaluate Eq. (21) with $R=1$ at a series of equally spaced time points $t=t_{i}$ from $t=0$ to $t=t_{\mathrm{obs}}$. When the transformation given in Eq. (24) subsequently requires the value of the solution to be evaluated at $t=t_{\mathrm{obs}}/R$, we approximate this using linear interpolation between the nearest two values of $t_{i}$.
@@
For the basic method, we first found the maximum likelihood estimate (MLE) for the target parameter set by solving the optimisation problem in Eq. (1). We then computed univariate likelihood profiles for each target parameter $\psi$ by taking a uniform mesh over some interval $[\psi_{\textrm{min}},\psi_{\textrm{max}}]$ with a modest number of mesh points ($n_{\mathrm{mesh}}=41$) and solving Eq. (6) to find the profile log-likelihood at each mesh point. In practice, we determined an appropriate interval $[\psi_{\textrm{min}},\psi_{\textrm{max}}]$ by computing $\hat{L}_{p}$ across a trial interval and if necessary widening the interval until it contained the 95% CI.
@@
We used a similar procedure for the structured method, finding the MLE by solving the nested optimisation problem in Eq. (4)–(5) (see Figure 1B). We then constructed univariate profiles as described above, but with Eq. (6) similarly recast as a nested pair of optimisation problems. This means that, with the structured method, there is one fewer parameter to profile because for a given dataset $x$, the inner parameter is treated as a deterministic function of the outer parameters and so does not need to be profiled separately.
Default parameter values used in model simulations for the three case studies are shown in Supplementary Table S1. All optimisation problems were solved numerically using the constrained optimisation function fmincon in Matlab R2022b with the interior point algorithm and default tolerances. For each model case study, we generated $M=500$ independent synthetic datasets by simulating the noise process on the solution of the relevant forward model. For each dataset, we calculated the relative error and the number of calls to the forward model required by both the basic and structured methods. This enabled us to compare the accuracy and efficiency of the two methods. We carried out sensitivity analysis by randomly varying model parameters for each synthetically generated dataset (see Supplementary Material for details).
Documented Matlab code that implements both the basic and the structured method on the three models described above is publicly available at: https://github.com/michaelplanknz/structured-inference. The code can also be run on a user-supplied model, either with synthetically generated data from the specified model or with user-supplied data. This requires the user to provide a function that returns the model solution for specified parameter values. The user can choose from a range of pre-supplied noise models, including additive or multiplicative Gaussian, Poisson, and negative binomial. Detailed instructions for running the method on a user-supplied models and/or data are available in the ReadMe file at the URL above.

### v3

@@ Abstract
Parameter inference and uncertainty quantification are important steps when relating mathematical models to real-world observations, and when estimating uncertainty in model predictions. However, methods for doing this can be computationally expensive, particularly when the number of unknown model parameters is large. The aim of this study is to develop and test an efficient profile likelihood-based method, which takes advantage of the structure of the mathematical model being used. We do this by identifying specific parameters that affect model output in a known way, such as a linear scaling. We illustrate the method by applying it to three toy models from different areas of the life sciences: (i) a predator-prey model from ecology; (ii) a compartment-based epidemic model from health sciences; and, (iii) an advection-diffusion-reaction model describing transport of dissolved solutes from environmental science. We show that the new method produces results of comparable accuracy to existing profile likelihood methods, but with substantially fewer evaluations of the forward model. We conclude that our method could provide a much more efficient approach to parameter inference for models where a structured approach is feasible. Code to apply the new method to user-supplied models and data is provided via a publicly accessible repository.
@@ Introduction
In 2014, Hines et al. [6] reviewed the application of Bayesian Markov chain Monte Carlo (MCMC) sampling methods for parameter estimation and parameter identifiability for a range of ordinary differential equation (ODE)-based models of chemical and biochemical networks. Their results indicated that parameter non-identifiability can be detected through MCMC chains failing to converge. Similar observations were made by Siekmann et al. [7] for a range of continuous-time Markov chain models used in the study of cardiac electrophysiology. In 2020, Simpson et al. [8] compared MCMC sampling with a profile likelihood-based method for parameter inference and parameter identifiability for a range of nonlinear partial differential equation (PDE)-based models used to interrogate cell biology experiments. The authors studied identifiable and non-identifiable problems and found that, while both the MCMC and profile likelihood-based approaches gave similar results for identifiable models, only the profile likelihood approach provided mechanistic insight for non-identifiable problems. They also found that profile likelihood was approximately an order of magnitude faster to run than MCMC sampling, regardless of whether the problem was identifiable or not. The speed-up in computation is a consequence of the fact that it is often faster to use numerical optimisation methods compared to sampling methods. Raue et al. [9] suggested the sequential use of profile likelihood to constrain prior distributions before applying MCMC sampling in the face of non-identifiability.
While profile likelihood has long been used to assess parameter identifiability and parameter estimation [10, 11, 12, 13, 14], Simpson and Maclaren [15] recently presented a profile likelihood-based workflow covering identifiability, estimation and prediction. The workflow uses computationally efficient optimisation-based methods to estimate the maximum likelihood, and then explores the curvature of the likelihood through a series of univariate profile likelihood functions that target one parameter at a time. This workflow then propagates uncertainty in parameter estimates into model predictions to provide insight into how variability in data leads to uncertainty in model predictions. While the initial presentation of the profile likelihood-based workflow focused on deterministic models, the same ideas can be applied to stochastic models whenever a surrogate likelihood function is available [1].
@@
As a motivating example, [16] described an epidemiological model that was fitted to data in real-time and used to provide policy advice during the Covid-19 pandemic. The model consisted of a system of several thousand ODEs, and model fitting and uncertainty quantification was done using an approximate Bayesian computation method targeting 11 unknown parameters. This resulted in a computationally intensive problem with a high-dimensional parameter space. Two of the fitted parameters were multiplicative factors on the proportion of model infections that led to hospitalisation and death respectively in each age group and susceptibility class. Adjusting either or both of these parameters while holding the other parameters fixed would linearly scale the time series for expected daily hospital admissions and deaths output by the model. Thus, for a given combination of the other nine fitted parameters, it should be possible to find the optimal values for these two parameters without a costly re-evaluation of the full model.
Here, we propose an new approach, which we term structured inference, that exploits the known scaling relationship between certain parameters and the model solution. Our approach is similar to that of Loos et al. [17], who developed a hierarchical parameter inference method for ODE models with additive Gaussian or Laplace noise where some of the parameters, which they termed ‘scaling parameters’, were simply multiplicative factors on expected value of the observed variables. This method was generalised by [18] to include offset parameters. Like [17, 18], our approach recasts an $N$-dimensional optimisation problem as a nested pair of lower-dimensional problems, effectively reducing the dimensionality of the search space. However, our methods go beyond those of [17, 18] in two important aspects: (1) they focus on parameter estimation, whereas we consider parameter estimation, practical identifiability and uncertainty quantification via the profile likelihood; (2) we design the method to be applicable in more general settings, including different classes of mathematical models than just ODE models, more general parameter relationships than just multiplicative scaling, and more general noise models.
We illustrate our method using canonical toy models representing three cases studies drawn from different areas of the life and physical sciences: ecological species interactions; epidemiological dynamics; and pollutant transport and deposition. These examples demonstrate some of the features described under (2) above, e.g. Poisson and negative binomial noise models in the first two examples, and a PDE model with a non-trivial parameter relationship in the third example. In each case study, we use the model structure to identify parameters that have a known scaling effect on model output. We then show that implementing a structured inference approach results in solutions with very similar accuracy, but substantially fewer calls to the forward model solver. This shows that our method could potentially offer significant reduction in computation time when applied to more complex models, which are computationally expensive to solve. Alongside this article, we provide fully documented code for implementing our method, with instructions for how it can be applied to user-supplied models and data.
@@ Methods
$y(\theta_{\mathrm{outer}},\theta_{\mathrm{inner}})=F\left(y(\theta_{\mathrm{outer}},\theta_{\mathrm{inner}}^{\mathrm{ref}}){;\theta_{\mathrm{inner}}}\right)$ (2)
for some known function $F$, which we assume is relatively cheap to compute compared to solving the full forward model for $y(\theta)$. The simplest example is where the model solution is directly proportional to $\theta_{\mathrm{inner}}$, in which case this transformation is a simple linear scaling:
@@
For the structured inference method, we use $p_{\mathrm{obs}}$ as the inner parameter $\theta_{\mathrm{inner}}$. We set the reference value of the inner parameter as $p_{\mathrm{obs}}^{\mathrm{ref}}=1$, noting that if ${\bf y}_{\mathrm{ref}}$ is the model solution with $p_{\mathrm{obs}}=1$, then the solution for any given value of $p_{\mathrm{obs}}\in[0,1]$ is simply ${\bf y}=p_{\mathrm{obs}}{\bf y}_{\mathrm{ref}}$. Thus, under the structured method, the ordinary differential equation model (7)–(8) only ever needs to be solved with $p_{\mathrm{obs}}=1$. When evaluation of the model likelihood requires the expected value ${\bf y}$ of observed data for any other value of $p_{\mathrm{obs}}$, this is obtained via this simple linear scaling. Note that the choice to set $p_{\mathrm{obs}}^{\mathrm{ref}}=1$ is arbitrary and any fixed value of $p_{\mathrm{obs}}^{\mathrm{ref}}\in(0,1]$ could be chosen.
SEIRS model: Epidemiology
Our second case study is a SEIRS compartment model for an epidemic in a closed population of size $N$ [24]. We assume that the population at time $t$ is divided into compartments representing the number of people who are susceptible ($S$), exposed ($E$), infectious ($I$), and recovered ($R$). To model waning immunity, we subdivide the recovered compartment into two compartments $R_{1}$ and $R_{2}$ and assume that individuals transition from $R_{1}$ to $R_{2}$ and subsequently back to $S$. This is described by the following system of ordinary differential equations:
@@
We assume that, on average, a fixed proportion $p_{\mathrm{obs}}$ of infections are observed. This could represent epidemiological surveillance data, for example the number of notified cases under the assumption that surveillance effort is steady over time. Alternatively it could represent a particular clinical outcome, such as hospital admission. Unknown case ascertainment or uncertain clinical severity are common, particularly for new or emerging infectious diseases. Hence, joint inference of this parameter with parameters of the SEIRS model, in situations where these are identifiable, is an important task in outbreak modelling [16, 25, 26].
@@
We assume that the number of observed cases $\hat{X}_{t}$ on day $t$ is drawn from a negative binomial distribution (which is a commonly used distribution in epidemic modelling when observed cases may have more stochasticity than under a binomial or Poisson distribution [27, 28]) with mean $y_{t}$ and dispersion factor $k$:
@@
Similar to the predator–prey model, we use $p_{\mathrm{obs}}$ as the inner parameter $\theta_{\mathrm{inner}}$ and set $p_{\mathrm{obs}}^{\mathrm{ref}}=1$, noting that if ${\bf y}_{\mathrm{ref}}$ is the model solution with $p_{\mathrm{obs}}=1$, then the solution for any given value of $p_{\mathrm{obs}}\in[0,1]$ is simply ${\bf y}=p_{\mathrm{obs}}{\bf y}_{\mathrm{ref}}$. Again, the choice to set $p_{\mathrm{obs}}^{\mathrm{ref}}=1$ is arbitrary and any fixed value of $p_{\mathrm{obs}}^{\mathrm{ref}}\in(0,1]$ could be chosen.
@@
$u(x,t)$ $=$ $\dfrac{U_{0}}{2}\left[\textrm{erfc}\left(\dfrac{x-vt/R}{\sqrt{4Dt/R}}\right)+\textrm{exp}\left(\dfrac{vx}{D}\right)\textrm{erfc}\left(\dfrac{x+vt/R}{\sqrt{4Dt/R}}\right)\right],$ (21)
@@
For the structured method, we exploit a symmetry in the analytical solution in Eq.(21), namely that
$u(x,t;R)=u\left(x,(R^{\mathrm{ref}}/R)t;R^{\mathrm{ref}}\right).$ (24)
We therefore use $R$ as the inner parameter $\theta_{\mathrm{inner}}$. We set $R^{\mathrm{ref}}=1$ and note that if $u^{\mathrm{ref}}$ is the model solution for $u(x,t)$ when $R=R^{\mathrm{ref}}=1$ then, by Eqs. (22) and (24), the solution for the product $s(x,t)$ for an arbitrary value of $R>1$ is
$s(x,t)=(R-1)u^{\mathrm{ref}}(x,t/R).$ (25)
This relationship between the inner parameter $R$ and the model solution is more complicated than the simple multiplicative scaling seen in the first two case studies. As Eq. (24) shows, changing the value of $R$ is equivalent to rescaling one of the independent variables ($t$) in the PDE for $u(x,t)$. Therefore, evaluating the solution $s(x,t)$ for a different value of $R$ is equivalent to evaluating the reference solution $u^{\mathrm{ref}}(x,t)$ at a different value of $t$, as well as changing the multiplicative factor that relates $u(x,t)$ and $s(x,t)$. Since a numerical scheme for an advection–diffusion PDE will typically generate solution values at a series of closely spaced time points, it will generally be straightforward to query the reference solution $u^{\mathrm{ref}}$ at an earlier time point. This allows the solution for arbitrary values of $R>1$ to be obtained without costly re-evaluation of the numerical solution scheme.
To emulate this situation, for given values of the outer parameters $D$ and $v$, we generate the reference solution $u^{\mathrm{ref}}$ by evaluating Eq. (21) with $R=1$ at a series of equally spaced time points $t=t_{i}$ from $t=0$ to $t=t_{\mathrm{obs}}$. When the solution for $s(x,t)$ in Eq. (25) subsequently requires $u^{\mathrm{ref}}$ to be queried at $t=t_{\mathrm{obs}}/R$, we approximate this using linear interpolation between the nearest two values of $t_{i}$. Here, choosing $R^{\mathrm{ref}}=1$ is beneficial because it means that the reference solution $u^{\mathrm{ref}}$ only ever needs to be queried at values of $t$ within the range of the numerical solution, $t\in[0,t_{\mathrm{obs}}]$.
@@
For each model case study, we first solved the underlying deterministic model equations with fixed parameter values (see Supplementary Table S1). This was done via numerical solution of the ODEs for the first two case studies, whereas the analytical solution was used for the third case study. Once the deterministic solution $y$ was calculated, we then generated a synthetic dataset $x$ by simulating the specified noise model.
To perform parameter inference using the basic method, we first found the maximum likelihood estimate (MLE) for the target parameter set by solving the optimisation problem in Eq. (1). We then computed univariate likelihood profiles for each target parameter $\psi$ by taking a uniform mesh over some interval $[\psi_{\textrm{min}},\psi_{\textrm{max}}]$ with a modest number of mesh points ($n_{\mathrm{mesh}}=41$) and solving Eq. (6) to find the profile log-likelihood at each mesh point. In practice, we determined an appropriate interval $[\psi_{\textrm{min}},\psi_{\textrm{max}}]$ by computing $\hat{L}_{p}$ across a trial interval and if necessary widening the interval until it contained the 95% CI.
@@
We used a similar procedure for the structured inference method, finding the MLE by solving the nested optimisation problem in Eq. (4)–(5) (see Figure 1B). We then constructed univariate profiles as described above, but with Eq. (6) similarly recast as a nested pair of optimisation problems. This means that, with the structured method, there is one fewer parameter to profile because for a given dataset $x$, the inner parameter is treated as a deterministic function of the outer parameters and so does not need to be profiled separately.
All optimisation problems were solved numerically using the constrained optimisation function fmincon in Matlab R2022b with the interior point algorithm and default tolerances. We calculated the relative error between the MLE and the true parameter values, and the number of calls to the forward model required by both the basic and structured methods. This enabled us to compare the accuracy and efficiency of the two methods. Because the results will depend on the particular realisation of the noise model, we repeated this process $M=500$ times for each case study, calculating the MLE and 95% CIs for $M$ independent datasets using fixed parameter values (see Supplementary Table S1). We report the median and interquartile range of the relative error and the number of function calls across the $M$ datasets. We also report the proportion of the $M$ datasets for which the 95% CI for each parameter contained the true value. We also carried out sensitivity analysis by randomly varying the model parameters for each synthetically generated dataset (see Supplementary Material for details).
Documented Matlab code that implements both the basic and the structured method on the three models described above is publicly available at: https://github.com/michaelplanknz/structured-inference. The code can also be run on a user-supplied model, either with synthetically generated data from the specified model or with user-supplied data. This requires the user to provide a function that returns the model solution for specified parameter values, and a separate function that transforms the reference model solution to the model solution for a specified value of the inner parameters. The user can choose from a range of pre-supplied noise models, including additive or multiplicative Gaussian, Poisson, and negative binomial. Detailed instructions for running the method on a user-supplied models and/or data are available in the ReadMe file at the URL above.